ER004B

PI: Ruscica
Title: BL Lacs

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 07:30(24/02)-02:00(25/02)
Stations: EbHhJbMcNtOnSmShTrUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

HartRAO (Jonathan Quick at Mon Feb 24 21:15:29 EST 1997) Minor failures
Read heads failing, write head status unknown

JIVE (E. Arbizzani at Mon Feb 24 21:38:59 EST 1997) Minor failures
Scan 1:24-1:37 - Total power integrator overflow IF2 - Tsys IF2 high No others problems know.

Medicina (E. Arbizzani at Mon Feb 24 21:39:28 EST 1997) Minor failures
Scan 1:24-1:37 - Total power integrator overflow IF2 - Tsys IF2 high No others problems know.

Westerbork (Tony Foley at Wed Feb 26 11:00:03 EST 1997) Minor failures
Removed 1 cryogenic frontend from tied-array for conversion to 18cm. 1 more telescope out because of pointing problems from 1700 to 1750 (high wind). The tape ran off at 2018 and was restarted at 2105UT.

Shanghai (Huang Xinyong at Wed Mar 5 05:12:11 EST 1997) Minor failures
Clock offset = 0.32 - 0.38 msec. First scan was missed because of receiver trouble. Weather overcast.

Jodrell (Alastair Gunn at Thu Mar 6 03:40:47 EST 1997) Failure
Did not observe due to high winds all day.

Calibration (Fredrik T Rantakyro at Thu May 8 11:31:00 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name er4b.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Mon Apr 6 10:57:01 METDST 1998) Minor failures
Correlation completed. Encountered the usual bad tracking for Torun. No other major problems.